OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (mysesshou):

f(x)=-4x+7 g(x)=10x-6 f(g(x)) we take the g(x) equation and input it into the x in the f(x) equation f(10x-6) = -4(10x-6)+7 then we solve for f(x) = -4(10x-6)+7 f(x) = -4(10x-6)+7 = -4(10x)-4(-6)+7 =-40x+24+7 =-40x+31